A baseball player tosses a baseball upward giving it an initial velocity of 8.20 m/s. How long does it take for the baseball to reach its maximum height? 


1
Expert's answer
2021-01-14T10:40:41-0500

We can find the time that the ball takes to reach its maximum height from the kinematic equation:


"v=v_0-gt,""0=v_0-gt,""t=\dfrac{v_0}{g}=\dfrac{8.20\ \dfrac{m}{s}}{9.8\ \dfrac{m}{s^2}}=0.84\ s."

Answer:

"t=0.84\ s."